OpenStudy (anonymous):

lim as x goes to 0+ In(sinx)

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0+}\ln(\sin(x))=\ln(0)=DNE\] DNE=Does not exist

OpenStudy (anonymous):

Pfft.. that's wrong It should be \[-\infty\]